OpenStudy (anonymous):

What is the five number summary for the following data set? 35,28,43,32,37,35,31,25,40,36

OpenStudy (amistre64):

how do you define a 5 number summary?

OpenStudy (anonymous):

its basically a box and whisker plot with a minimum , maximum etc

OpenStudy (amistre64):

yeah, why they didnt say box and whisker .... first sort the numbers; thatll tell you min and max, and give us a way to pull the rest

OpenStudy (amistre64):

there are 10 numbers, so the median is in the middle of the 2 center values

OpenStudy (amistre64):

finding quartiles is just dividing the number of elements by 4 to determine the positional place of q 10/4 = 2.5; Q1 = 1(2.5) rnded up Q2 = 2(2.5) is a hole number so average is required Q3 = 3(2.5) rnded up

OpenStudy (amistre64):

so the hardest part really is sorting the numbers ....

OpenStudy (anonymous):

the numbers in order are 25,28,31,31,35,36,40,43

OpenStudy (amistre64):

25, 28, 31, 31, 35, 36, 40, 43 ... thats only 8 numbers, your missing 2

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sort%7B35%2C28%2C43%2C32%2C37%2C35%2C31%2C25%2C40%2C36%7D 25, 28, 31, 32, 35, 35, 36, 37, 40, 43 0 1 2 3 4 avg=35

OpenStudy (anonymous):

so the maximum is 43 , the minimum is 25